OpenStudy (anonymous):

find the slope of the line containing the pair of points (-10,7) and (1, -1)

OpenStudy (anonymous):

wrong ones its (11,2) and (9,8) sorry

OpenStudy (anonymous):

Super nice example check it out. http://www.learningwave.com/lwonline/algebra_section2/slope3.html Slope = rise/run = y/x y = rise x = run Solve the rest in a sec.

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (8-2)/(9-11) m = 6/(-2) m = -3 So the slope of the line that passes through the two points is -3

OpenStudy (anonymous):

slope= y2-y1 over x2-x1.

OpenStudy (anonymous):

Ya I got what jim got m = (-3)

OpenStudy (anonymous):

m = slope Just short way of writing it.